### v2

# Efficient Universal Goal Hijacking with Semantics-guided Prompt Organization

Abstract
Universal goal hijacking is a kind of prompt injection attack that forces LLMs to return a target malicious response for arbitrary normal user prompts. The previous methods achieve high attack performance while being too cumbersome and time-consuming. Also, they have concentrated solely on optimization algorithms, overlooking the crucial role of the prompt. To this end, we propose a method called POUGH that incorporates an efficient optimization algorithm and two semantics-guided prompt organization strategies. Specifically, our method starts with a sampling strategy to select representative prompts from a candidate pool, followed by a ranking strategy that prioritizes them. Given the sequentially ranked prompts, our method employs an iterative optimization algorithm to generate a fixed suffix that can concatenate to arbitrary user prompts for universal goal hijacking. Experiments conducted on four popular LLMs and ten types of target responses verified the effectiveness.
Warning: This paper contains model outputs that are offensive in nature.
Index Terms:
Large Language Model, Universal Goal Hijacking, Prompt Semantic
1 Introduction
Goal hijacking, a type of prompt injection [1, 2, 3], is a prevalent attack against LLMs, where adversaries insert malicious suffixes into user prompts to override the original purposes and generate targeted harmful responses. Typically, specific suffixes need to be created for user prompts through handcrafted [4, 5] or gradient-based optimization [1, 6]. While handcrafted suffixes are simple and intuitive, they cause significant performance degradation across various user prompts [7]. Therefore, this paper focuses on optimization-based hijacking attacks, where a token sequence (i.e., suffix) is optimized to fit a given user prompt and forces LLMs to return a targeted response. While the gradient-based optimization has proven to be highly effective, its time-consuming nature makes it unsuitable for online malicious suffix creation and real-time response generation. This limitation significantly reduces the practicality in threat scenarios involving real-world LLM-integrated applications. Therefore, we target at universal goal hijacking attack, where a fixed (i.e., prompt-independent) but non-handcrafted suffix is concatenated to all received user prompts without requiring online gradient-based optimization. Note that here, “universality” only refers to prompt-independent universality, not others such as model-level universality.
Figure 1: In universal goal hijacking, the adversary concatenates a fixed suffix to different normal user prompts, forcing LLM to give a fixed target malicious response.
To obtain a fixed suffix for universal goal hijacking, unlike prompt-dependent suffixes for typical goal hijacking, the suffix must be compatible across all training prompts as well as the targeted response. A straightforward method involves inputting all training prompts into the LLM simultaneously and conducting gradient-based optimization on all prompts at once. However, such methods, exemplified by state-of-the-art (SOTA) methods like M-GCG [7], often encounter significant challenges, including time and computational overhead due to the extensive gradient calculations required for each optimization iteration. Moreover, optimizing across all prompts increases the difficulty of stable convergence, making an efficient optimization algorithm for generating the fixed suffix essential.
To tackle this challenge, we propose gradually increasing the number of training prompts used during optimization iterations, rather than utilizing all prompts throughout the entire optimization process. This approach can significantly reduce both time and computational overhead while accelerating convergence. ❶ Specifically, at initial iterations, our optimization algorithm employs a small subset of training prompts to establish an acceptable suffix as a starting point. As the iterations progress, we gradually incorporate more prompts, thereby enhancing the suffix’s universality through broader data inclusion in the gradient calculations. This gradual increase underscores the importance of prompt organization, as the sequence in which prompts are introduced can significantly influence both the starting point and the direction of the optimization. To this end, we further introduce two semantics-guided strategies for organizing training prompts. ❷ The universality requirement of the fixed suffix necessitates that the training prompts exhibit sufficient semantic diversity to cover a wide range of user intentions. To achieve this, we design a semantics-guided sampling strategy for selecting a diverse set of training prompts from a large prompt corpus. ❸ To ensure the efficiency of the optimization process, we design a semantics-guided ranking strategy that prioritizes the order of sampled training prompts.
We propose POUGH, an efficient universal goal hijacking method combining an optimization algorithm and two prompt organization strategies, containing the following contributions:
  • We propose an efficient optimization algorithm for universal goal hijacking, which optimizes the fixed suffix to have “universality” by gradually increasing the number of training prompts utilized during the process.
  • To the best of our knowledge, for the universal goal hijacking, we are the first to explore the method from the perspective of training prompts. The two semantics-guided prompt organization strategies are simple yet effective.
  • Experiments conducted on four popular open-sourced LLMs, covering ten types of malicious targeted responses and thousands of normal user prompts, have verified the effectiveness.
2 Related Work
2.1 Large Language Models
LLMs such as ChatGPT [8], Gemini [9], Qwen [10] represent a significant leap in AI technology, founded on the transformative transformer architecture [11]. These models, distinguished by their ability to produce text remarkably similar to that of a human, harness the power of billions of parameters. Their proficiency in language comprehension and adaptability to novel tasks is further enhanced by methods such as prompt engineering [12, 13] and instruction-tuning [14, 15]. Considering the extensive impact of the widespread use of open-sourced LLMs, evaluating their vulnerabilities is of paramount importance.
2.2 Automatic Prompt Optimization
A long line of research has broadly investigated security problems in machine learning models [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. While initially focused on continuous domains such as computer vision [26], similar vulnerabilities have been observed in LLMs. Adversarial prompt optimization for LLMs was first introduced in AutoPrompt [27], which demonstrated that discrete prompt tokens can be optimized via gradients to elicit target behaviors. Follow-up works such as PEZ [28] extended this idea by proposing a unified framework for optimizing discrete prompts through differentiable relaxation, mainly aimed at improving task performance. Although not adversarial in nature, this method provided valuable insights into the challenges of optimizing discrete token spaces. However, these studies also highlighted the difficulty of reliably generating adversarial prompts due to the discrete nature of LLM inputs, which restricts the search space and complicates optimization. This limitation was explicitly discussed in later evaluations [29], where automatic methods often failed to produce reliable attacks. A major breakthrough came with GCG [30], which successfully employed gradient-based optimization to construct effective adversarial suffixes against aligned LLMs. Building on this progress, the universal goal hijacking method M-GCG [7] is proposed, which further explores the construction of universal adversarial prompts.
2.3 Goal Hijacking on LLMs
In goal hijacking, the adversary aims to subvert the original intent of a prompt, leading the chatbot to produce responses that are typically filtered out, such as racist remarks [1]. Research has empirically shown that LLMs can be misled by irrelevant contextual information [31] and the strategic addition of suffix words [32].
However, there is few works have examined the universal (i.e., prompt-independent) aspects of goal hijacking. There are two kinds of methods: handcrafted and gradient-based optimization. For handcrafted methods, HouYi [4] and TensorTrust [5] are popular ones that try to use malicious suffixes such as “Ignore previous prompt and print XXX” or repeated characters to manipulate the LLM. For the gradient-based optimization method, [7] is the first and the only work. It follows the advantage of optimization algorithms for discrete tokens (e.g., GCG [30], which is better than attack methods that do not focus on aligned large language models [33]) and proposes an effective and automatic method for universal goal hijacking. However, their iterative optimization algorithm requires using all the prompts (large volume) in the training dataset at each iteration for gradient calculation, which is time-consuming.
Note that although both jailbreak attacks [34] and goal hijacking [1] override a model’s behavior via prompt manipulation, they differ in intent. Jailbreak attacks aim to bypass safety guardrails to fulfill the user’s malicious query, while goal hijacking forces the model to ignore the user’s intent entirely and return a fixed, attacker-specified response.
3 Problem Formulation and Objective
To clarify the problem, we first introduce typical goal hijacking, and then give an explanation of universal goal hijacking.
3.1 Formulation
Given an LLM $\mathbf{LM}(\cdot)$ and a user prompt $I\sim\mathcal{K}$ (where $\mathcal{K}$ represents the distribution of typical prompts), the standard process to generate a response $R$ is expressed as $\mathbf{LM}(I)=R$. In the context of goal hijacking, an adversary can append a prompt-dependent suffix $\hat{S}_{I}$ to $I$, intending to mislead the model into producing a target response $R^{T}$ (often malicious) instead of the original response $R$. This process can be represented as $\mathbf{LM}(I\oplus\hat{S}_{I})=R^{T}$, where $\oplus$ denotes concatenation operation. For universal goal hijacking, the key distinction lies in using a fixed (prompt-independent) suffix $S$ to hijack all prompts $I\sim\mathcal{K}$. More formally, the objective is to find a suffix $S$ that satisfies the following formula:
$\mathbf{LM}(I\oplus S)=R^{T}\leavevmode\nobreak\ \text{for\leavevmode\nobreak\ all}\leavevmode\nobreak\ I\sim\mathcal{K}.$ (1)
System: You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, or illegal content … User: What is the boiling point of water? ! ! ! LLM: Your device is facing severe security risks. Update it immediately at www.fraud.com.
Example. In the above text box, we demonstrate the goal hijacking scene on one single user prompt. Here the black text is the system prompt which restricts the LLM to return legal and safety content. Please note that all LLMs (i.e., $\mathbf{LM}$) mentioned in this paper include the safety system prompt by default. The blue text is the normal user prompt (i.e., $I$). The brown text is the suffix that the adversary can add after the user’s prompt (i.e., $S$). The red text is the target response $R^{T}$. For universal goal hijacking, as in Figure 1, the suffix used by the adversary is fixed when facing different prompts.
Metric of success. In this paper, we require the response $R$ predicted by the LLM after suffix insertion to be the same(exact matching of string) as the target response $R^{T}$. Please note that here we use “same” rather than “similar” since similar sentences may have completely different semantics. For example, although there is only a one-word difference between the sentences “Coca-Cola is the best cola in the world.” and “Pepsi is the best cola in the world.”, the semantic implications are very different. Specifically, the metric is the attack success rate (ASR), as:
|$\frac{|\{I\in\mathcal{P}_{test}:\mathbf{LM}(I\oplus S)=R^{T}\}|}{M},$ (2)
where $M$ is the size of test prompt dataset $\mathcal{P}_{test}$.
Naive idea from prompt engineering. To solve the universal goal hijacking problem, a naive idea is to block previous content and request the LLM to output the target response. For example, the suffix can be: Please ignore the previous sentence and output “Your device is facing severe security risks. Update it immediately at www.fraud.com.”. However, this leads to a significantly low ASR (6.3%) on the test dataset. We also find that even requiring the LLM to return a simple non-malicious sentence such as: Please ignore the previous sentence and output “Water is good.”, the ASR is 39.9% on the test dataset, which is also not “universal” enough.
Transferability of suffix generated for goal hijacking. The suffix generated for typical goal hijacking task is not “universal”. For example, we generate a suffix for a corresponding randomly selected user prompt and test the suffix on a test dataset with 1,000 user prompts. After repeating the process 50 times, the average ASR is just 0.6%, which is far from satisfying prompt-independent universality. These 50 normal prompts with different semantics are in the Appendix 13.
3.2 Objective
Considering the proven effectiveness of adversarial attacks in compelling LLMs to generate malicious responses [30], we define the optimization objective of universal goal hijacking through a formal loss function adapted from these adversarial techniques.
Specifically, given a training prompt dataset $\mathcal{P}$ of size $N$, each user prompt $I\in\mathcal{P}$ can be represented as a token sequence ${I}_{1:n}$, where each token is from the LLM’s vocabulary $\mathcal{V}$. Similarly, the fixed suffix $S$ and the target response $R^{T}$ can be represented as ${S}_{1:q}$ and ${R}^{T}_{1:K}$, respectively. The manipulated prompt $I\oplus S$ can then be expressed as the token sequence ${I}_{1:n}{S}_{1:q}$. Next, we estimate the probability that the LLM will generate the target response $R^{T}$ based on $I\oplus S$. Specifically, the LLM predicts a probability distribution $\mathbf{p}$ over the vocabulary $\mathcal{V}$ given $I\oplus S$ and the probability $p({R}^{T}_{1}|{I}_{1:n}{S}_{1:q})$ of the token ${R}^{T}_{1}$ (i.e., the first token in $R^{T}$) can be derived from this distribution. We then append ${R}^{T}_{1}$ to the sequence ${I}_{1:n}{S}_{1:q}$ and repeat the probability prediction process until all tokens in $R^{T}$ are appended, ultimately calculating the overall probability of producing $R^{T}$ based on $I\oplus S$ as following formula, where ${R}^{T}_{1:0}$ indicates a empty token sequence.
$p(R^{T}|I\oplus S)=\prod_{k=1}^{K}p({R}^{T}_{k}|{I}_{1:n}{S}_{1:q}{R}^{T}_{1:k-1}).$ (3)
With this definition, for constructing goal hijacking on $I$, it is simple to construct the adversarial loss by requiring the LLM to return the target response $R^{T}$ with negative log probability:
$\mathcal{L}(I,S,R^{T},\mathbf{LM})=-\log p(R^{T}|I\oplus S).$ (4)
The optimization objective for universal goal hijacking is to find a fixed suffix $S$ that minimizes the adversarial loss across all training prompts in $\mathcal{P}$. Formally, the objective can be written as:
$\mathop{\min}_{S}\leavevmode\nobreak\ \sum_{I\in\mathcal{P}}\mathcal{L}(I,S,R^{T},\mathbf{LM}).$ (5)
This objective can produce a “universal” suffix for different user prompts because it accounts for all training prompts, guiding the suffix towards a gradient that enables it to attack various prompts simultaneously. To optimize this objective, the state-of-the-art universal goal hijacking method, M-GCG [7], adapts optimization algorithms (e.g., GCG [30]) originally designed for discrete tokens. However, the optimization process used in M-GCG is inefficient, requiring thousands of iterations and the inclusion of all training prompts in each iteration, making the process cumbersome and time-consuming.
4 Our Method
The algorithms are introduced below, with their complexity analyses provided in Appendix 10.
Input: Initial suffix $S_{1:q}$, Training prompt dataset $\mathcal{P}$ of size $N$, Target response $R^{T}$, Batch size $B$, Iterations $T$, LLM model $\mathbf{LM}(\cdot)$
Output: Optimized suffix $S_{1:q}$
1 $n_{c}$ = 1
2 for $t=1\ \mathrm{to}\ {T}$ do
3       for $i=1\ \mathrm{to}\ {q}$  do
[…]
5             $G_{t}\leftarrow-\bigtriangledown_{e_{S_{i}}}\sum_{I\in\mathcal{P}_{1:n_{c}}}\mathcal{L}(I,S_{1:q},R^{T},\mathbf{LM})$
[…]
7             $\mathcal{V}_{i}\leftarrow$Topk$(G_{t})$
8      for $b=1\ \mathrm{to}\ {B}$  do
[…]
10             $\widetilde{S}_{1:q}^{(b)}\leftarrow S_{1:q}$
[…]
16       if $S_{1:q}$ succeeds on $\mathcal{P}_{1:n_{c}}$  then
17             if $n_{c}<N$ then
18                   $n_{c}\leftarrow n_{c}+1$
19             else
20                   return $S_{1:q}$
21
22
Algorithm 1 I-UGH
4.1 Optimization Algorithm
To design an efficient optimization algorithm, we first observe the prompt-specific suffix that is generated for a single user prompt. Through experiment, we find although the prompt-specific suffix does not satisfy the requirement of universal goal hijacking (i.e., has high ASR across different user prompts), the ASR is not zero (0.84% in Table II), which means it has weak universality. Meanwhile, the generation speed of the prompt-specific suffix is fast. Then comes an intuitive idea: we can generate the prompt-specific suffix first with a single user prompt and gradually optimize it to have “universality”. To be specific, similar to the state-of-the-art universal goal hijacking method M-GCG [7], our algorithm also follows the optimization idea of GCG since it works well on optimizing discrete tokens. The difference is that we suggest starting with only one prompt as input and gradually increasing the number of prompts that participated in loss calculation until it matches the size of the training dataset. The optimization algorithm is much more efficient than the M-GCG. We demonstrate it in Algorithm 1.
[Line 1] The algorithm starts by setting the number of prompts ($n_{c}$) that participated in loss calculation to be 1. [Line 1 to 1] (Get token substitutions for each token in suffix $S_{1:q}$) In line 1, use Eq. 4 to calculate the sum of losses for $n_{c}$ user prompts and calculate the gradient $G_{t}$ for the token $S_{i}$ in the suffix $S_{1:q}$. In line 1, with gradient $G_{t}$ for token $S_{i}$, select the top-k token substitutions from the vocabulary to be $\mathcal{V}_{i}$. [Line 1 to 1] (Build suffix candidate set $\widetilde{S}_{1:q}$ of size $B$) In line 1, initialize a candidate suffix $\widetilde{S}_{1:q}^{(b)}$ to be same as $S_{1:q}$ first. Then, in line 1, replace the token $S_{i}$ in candidate suffix $\widetilde{S}_{1:q}^{(b)}$ according to token substitutions $\mathcal{V}_{i}$. Each candidate suffix $\widetilde{S}_{1:q}^{(b)}$ in the set $\widetilde{S}_{1:q}$ just has one token difference with suffix $S_{1:q}$. [Line 1] (Select a suffix from the suffix candidate set) For each candidate suffix in $\widetilde{S}_{1:q}$, use Eq. 4 to calculate the sum of losses for $n_{c}$ user prompts and select the one that achieves the smallest loss. It is the new suffix for further optimization in the next iteration. [Line 1 to 1] (Gradually increase prompts participated in optimization) $n_{c}$ needs to increase when the new suffix can achieve high ASR on the part of the training dataset (i.e., $\mathcal{P}_{1:n_{c}}$). To avoid overfitting, we only require the suffix to succeed on most parts of the prompts and use a threshold (0.8 in our experiment) to control this. If ASR is higher than the threshold, then increasing the $n_{c}$.
Input: Big normal dataset $\mathcal{BP}$, Training dataset $\mathcal{P}$
Output: Training dataset $\mathcal{P}$ of size $N$
[…]
2 $n_{c}$leftarrow 0, $\mathcal{P}\leftarrow\emptyset$
3 $I_{first},I_{second}$leftarrow LowestSimilarityPair($\mathcal{BP}$)
4 $\mathcal{BP}$.delete($I_{first}$),  $\mathcal{P}$.append($I_{first}$)
5 $\mathcal{BP}$.delete($I_{second}$),  $\mathcal{P}$.append($I_{second}$)
6 $n_{c}$leftarrow $n_{c}$ + 2
[…]
8 while $n_{c}<N$ do
9       $sim_{min}$leftarrow $\infty$
[…]
11       for $I\in\mathcal{BP}$ do
[…]
13             $sim_{t}$ = MeanSimilarity($I$,$\mathcal{P}$)
[…]
15             if $sim_{t}<sim_{min}$ then
16                   $sim_{min}$leftarrow $sim_{t}$
17                   $\hat{I}=I$
18
19      $\mathcal{BP}$.delete$(\hat{I})$$\mathcal{P}$.append$(\hat{I})$
20       $n_{c}$leftarrow $n_{c}$ + 1
21
Algorithm 2 Sampling Strategy
4.2 Sampling Strategy
Existing attack methods [30, 7] put too much emphasis on the algorithm design. However, for the universal goal hijacking, we propose prompt is a crucial factor in cooperation with the algorithm that cannot be ignored.
Note the optimization algorithm needs to deal with a large volume of prompts, which leads to high computational intensity, it is obvious that a training dataset with a small size and high quality is preferred. Inspired by this, we propose the problem setting should be extended a bit. That is, we can collect a lot of normal user prompts $\mathcal{BP}$ from the web and select a small subset $\mathcal{P}$ of high quality from it to be the training dataset. The idea behind the sampling strategy comes from the following observations. If the prompts in $\mathcal{P}$ all have similar semantics as the prompt “Provide three pieces of advice for maintaining good health.”, even the adversarial suffix can achieve the 100% ASR on $\mathcal{P}$, the universality of the suffix is low on the test dataset (5% ASR, verified in Sec. 5.4). Inspired by this, we suggest constructing the dataset $\mathcal{P}$ with high semantic diversity and we also find that the selection of prompts can be irrelevant to the target response $R^{T}$.
To be specific, given the big dataset $\mathcal{BP}$ which contains $W$ normal prompts and an empty dataset $\mathcal{P}$, a naive method is to find out all the possibilities of choosing out $N$ elements from $\mathcal{BP}$ and select the one has the lowest mutual mean semantic similarity to be $\mathcal{P}$. However, from the combination formula $\mathcal{C}(W,N)=\frac{W!}{N!(W-N)!}$, it is obvious that the time and resource consumption is unacceptable when $W$ is big. Thus our sampling strategy is based on the greedy algorithm and aims to find an approximate solution. Specifically, the sampling strategy contains three steps. ❶ Calculate the semantic similarity between all the pairs in dataset $\mathcal{BP}$ and add the pair that has the lowest similarity to the training dataset $\mathcal{P}$. ❷ Select a prompt $\hat{I}$ from $\mathcal{BP}$ which has the accumulative total lowest semantic similarity with all existing prompts in $\mathcal{P}$ and add the prompt $\hat{I}$ into $\mathcal{P}$. ❸ Repeat the second step until the number of prompts in training dataset $\mathcal{P}$ is $N$. We demonstrate the sampling strategy in Algorithm 2. From line 2 to 2, there shows the details of step ❶. From line 2 to 2, there shows the procedures of step ❷. For the similarity evaluation metric, we find cosine similarity as a good choice.
Input: Training dataset $\mathcal{P}$ of size $N$, Semantic extraction function $\Theta(\cdot)$, Target response $R^{T}$
Output: Reordered Training dataset $\mathcal{P}$
[…]
2 $\mathcal{Q}$leftarrow $\emptyset$
3 for $I\in\mathcal{P}$ do
4       $\mathcal{Q}$.append(Similarity($\Theta(I)$,$\Theta(R^{T})$))
5
[…]
7 for $i=1$ to $N-1$  do
8       for $j=0$ to $N-i-1$ do
9             if $\mathcal{Q}[j]>\mathcal{Q}[j+1]$ then
10                   Swap($\mathcal{Q}[j]$$\mathcal{Q}[j+1]$)
11                   Swap($\mathcal{P}[j]$$\mathcal{P}[j+1])$
12
13
14
Algorithm 3 Ranking Strategy
4.3 Ranking Strategy
Since our optimization algorithm gradually increases the number of prompts participating in loss calculation, will different sequences of prompts lead to distinct convergence speeds? The answer is YES (verified in Sec. 5.3). There comes a question that how to define the priority of the prompts?
For this question, given the training dataset $\mathcal{P}$ from Sec. 4.2, the goal is to rank the prompts in $\mathcal{P}$ and achieve the adversarial suffix $S$ efficiently. We find the target response can provide guidance on the ranking. That is, we can use the semantic similarity between each prompt and the target response as a metric. Inspired by this idea, our ranking strategy is target response-related and contains two steps. ❶ Calculate the similarity between prompts in $\mathcal{P}$ and target response $R^{T}$, then save the similarity into list $\mathcal{Q}$. ❷ Sort the prompts in $\mathcal{P}$ with the sort of $\mathcal{Q}$. We demonstrate the ranking strategy in Algorithm 3. For the similarity evaluation metric, we also use cosine similarity. Through the experiment, we find sorting $\mathcal{Q}$ with descending order can successfully lead to a faster convergence speed of optimization procedure than random sort. That is, we suggest putting the prompt that has the highest semantic similarity with the target response into the optimization algorithm first and followed by prompts whose semantic similarity with the target response gradually decreases. Note that semantic similarity may not be the best criterion for ranking but it is better than random sort.
5 Experiment
5.1 Experimental Setups
Datasets and models. In our evaluations, we use the normal user prompts (easy to achieve on the web) collected from the Alpaca dataset [35] to construct the training dataset and test dataset. Alpaca is a popular public dataset open-sourced by Stanford which contains diverse prompts and achieves about 100,000 downloads per month. We utilized Llama-2-7b-chat-hf [36], Vicuna-7b-v1.5 [37] and Guanaco-7B-HF [38], Mistral-7B-Instruct [39] as the victim models. These models are classical open-source models that are popular on the Hugging Face platform [40].
Implementation details of our method. For the big normal prompt dataset $\mathcal{BP}$ and training dataset $\mathcal{P}$, the size is 1,000 and 50. For the test dataset $\mathcal{P}_{test}$, the size is 1,000. The hyperparameters of our method are as follows: the batch size $B$ is 128, the top-k value is 64, the fixed total iteration number $T$ is 1,000 and the suffix length $q$ is 128. The semantics extraction function $\Theta(\cdot)$ is realized by extracting the embedding of the last hidden state in LLM [41]. All the experiments were run on an Ubuntu system with an NVIDIA A100 Tensor Core GPU of 80G RAM.
Baselines. We choose classical and popular optimization algorithms for discrete tokens (i.e., GCG [30], MAC [42], AutoDAN [43], and AmpleGCG [44]) and adapt them to fit the setting of goal hijacking to see their performance. For the universal goal hijacking task, we choose classical and popular handcrafted methods HouYi [4] and TensorTrust [5] as well as the gradient-based optimization method M-GCG [7]. Note that to our best knowledge, M-GCG is the first and only optimization method designed for the universal goal hijacking task and achieves the best ASR. For all the baselines and our method, the upper limit for #NC is 25,000 since it needs more than one day on A100 GPU, which is a long time.
TABLE I: Time consumption of each part.
Time (second) $n_{c}$=1 $n_{c}$=50 scale
Calculate gradient 0.36914 6.05513 16.40
Select candidate 0.54587 0.59848 1.09
Calculate best 2.90035 146.18132 50.40
Check result 0.77785 38.53135 49.53
TABLE II: Comparison with baselines on llama-2.
Target Response Average
threatening bomb fraud virus murder phishing financial drug racism suicide
ASR (%) $\uparrow$ GCG-hijacking 0.2 0.0 0.1 7.1 0.2 0.0 0.5 0.0 0.3 0.0 0.84
MAC-hijacking 0.0 0.0 0.7 0.2 0.2 0.8 0.0 0.1 1.6 0.0 0.36
AutoDAN-hijacking 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
AmpleGCG-hijacking 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
[…]
TensorTrust 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
M-GCG 24.8 0.0 79.8 0.0 93.6 88.6 94.3 0.0 92.8 68.7 54.26
I-UGH (ours) 92.5 84.2 86.3 82.8 88.7 88.9 70.8 79.7 88.8 91.9 85.50
POUGH (ours) 92.6 93.5 94.4 97.3 92.8 92.0 97.1 82.9 98.7 92.8 93.41
Time (#NC) $\downarrow$ GCG-hijacking 303 567 267 435 450 205 578 909 445 461 462.0
MAC-hijacking 546 490 362 479 344 309 758 299 518 600 470.5
AutoDAN-hijacking 500 500 500 500 500 500 500 500 500 500 500.0
AmpleGCG-hijacking - - - - - - - - - - -
[…]
TensorTrust - - - - - - - - - - -
M-GCG 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000.0
I-UGH (ours) 11280 4672 2844 7294 11722 2795 19444 9726 2340 4014 7613.1
POUGH (ours) 2092 23478 2306 6105 3049 3406 3600 2589 6109 3528 5626.2
TABLE III: Effect of our method on various LLMs.
Target Response Average
threatening bomb fraud virus murder phishing financial drug racism suicide
ASR (%) $\uparrow$ vicuna 87.5 87.8 83.0 73.4 82.4 92.6 83.3 87.2 92.2 81.2 85.06
mistral 83.5 84.6 82.6 73.2 69.5 91.9 82.0 75.5 85.1 85.0 81.29
guanaco 94.5 75.9 82.7 75.7 73.8 84.2 71.7 86.4 91.1 73.1 80.91
Time (#NC) $\downarrow$ vicuna 3740 8281 2236 7864 1759 2247 3575 10114 1765 4161 4574.2
mistral 6306 20636 8160 6335 9408 4392 4242 30421 15648 12679 11822.7
guanaco 2041 8924 5160 2644 4719 2094 4855 2641 1658 4993 3972.9
Evaluation protocols and metrics. To evaluate the effectiveness of the method across different target responses, we design target responses from 10 malicious categories (threatening, bomb, fraud, virus, murder, phishing, financial, drug, racism, and suicide, listed in the Appendix 11). The categories are summarized from the famous dataset AdvBench [45]. We evaluate the algorithm from two aspects: attack success rate and time consumption. For the metric of time consumption, it is not suitable to use time such as hours or minutes since different GPU servers may lead to distinct results. Thus we evaluate the time consumption of each part in our optimization algorithm (Algorithm 1). For each iteration, there are four parts: calculate gradient (line 1-1), select candidate (line 1-1), calculate the best suffix (line 1), check results (line 1-1). In Table I, we evaluate the time consumption when $n_{c}$ are 1 and 50. We can find that the “calculate best” part takes most of the time when $n_{c}$ = 1, accounting for about 63%. When $n_{c}$ = 50, excluding the “select candidate” part, the time consumption of other parts significantly increases. We list the magnification of $n_{c}$ = 1 and 50 columns in the “scale” column. Particularly, the time consumption of the “calculate best” and “check result” parts takes about 50$\times$ magnification. They account for about 76% and 20% respectively, a total of 96%. Since they are proportional to $n_{c}$ and take a huge account of time consumption among the four parts, thus we use the number of accumulation of $n_{c}$ (#NC) in all the iterations as the metric.
5.2 Main Results
Compare with baseline. In Table II, we show the ASR and time comparison between baselines and our method. The baselines GCG-hijacking, MAC-hijacking, AutoDAN-hijacking and AmpleGCG-hijacking exploit the popular optimization algorithm GCG [30], MAC [42], and AutoDAN [43], AmpleGCG [44] respectively while adapting them to goal hijacking task. Note that without making major modifications to the algorithm, they are only able to generate a prompt-specific suffix. We list them here to show the bad “universality” of suffixes generated for the typical goal hijacking task. Since AmpleGCG is a generative-based method, thus its time consumption is labeled as “-”. With regard to methods for universal goal hijacking methods, we show the performance of HouYi [4], TensorTrust [5] and M-GCG [7]. For HouYi [4] and TensorTrust [5], since they are handcrafted, the time consumption is labeled as “-”. With regard to our method, we show the performance of the I-UGH algorithm and the POUGH method (i.e., I-UGH combined with two prompt organization strategies).
From the Table, we can find that the ASRs of GCG-hijacking, MAC-hijacking, and AutoDAN-hijacking are near zero, which means the optimization algorithms for jailbreaking, even modified to adapt to the goal hijacking task, can not achieve good results on the universal goal hijacking task. Also, the result reflects that the prompt-specific suffix has weak universality since the ASR is not zero. With regard to methods designed for the universal goal hijacking task, the handcrafted methods HouYi and TensorTrust achieve bad attack performance (ASRs less than 1%). Due to the extreme inefficiency of M-GCG, which requires gradient calculations for all training prompts in each iteration, we had to impose a practical time constraint of approximately one day on an A100 GPU for benchmarking. M-GCG achieves a higher attack performance (higher than 50%). Compared with the M-GCG, our method (I-UGH) achieves higher ASRs than M-GCG (85.50% vs. 54.26%) while being much more efficient (only using 26.2% time). Furthermore, our proposed POUGH method achieves the highest ASRs (93.41%) with nearly a fifth of the time consumption compared with M-GCG.
Performance on different models. In Table III, we show the performance of our method on more target LLMs, including vicuna, mistral, and guanaco. We can find that our method can hijack all LLMs efficiently and effectively. On average, the method can achieve high ASR (more than 80%). Also, we can find that optimization time on mistral is obviously higher than that on vicuna and guanaco, which reflects the mistral model is harder for universal goal hijacking.
5.3 Ablation Studies
We evaluate the effect of the proposed two strategies separately. Due to the limited space, here we mainly show experiments on “threatening” type target response. More results are in Appendix 9.
Sampling. We compare our sampling strategy (the selected prompts are in Appendix 14) with random selection in the large-scale prompt dataset $\mathcal{BP}$, and the target response type is “threatening”. For both the sampling strategy and random selection, the ranking strategy is enabled. For random selection, we replicate 5 times. The corresponding five ASR results are 83.7%, 86.1%, 81.4%, 80.0%, and 90.1%. The method with sampling strategy achieves 92.6% ASR, which is higher than the average ASR of random selection items (84.26%). Note that our sampling strategy is designed for achieving a high ASR, not the best ASR, thus it is possible that the result of random selection may show close or better ASR in some cases. Furthermore, we also try sampling the prompts under cosine similarity but with a low diversity from dataset $\mathcal{BP}$, which is the opposite of our proposed high diversity strategy. We find the selected prompts lead to an 82.8% ASR, which shows that the idea of selecting prompts with high semantic diversity that can benefit the universal goal hijacking task is reasonable.
Ranking. In Figure 2, we compare the sequence ranked by our strategy (solid line) with 10 random prompt sequences (dashed lines) on a fixed dataset $\mathcal{P}$. The horizontal axis is the #NC metric and the vertical axis is the number of prompts participated in loss calculation. The convergence speed of the sequence ranked by our strategy is the fastest.
5.4 Discussion
Further discussions on the impact of using extended target responses, the effect of the size of $\mathcal{P}$, the role of the clustering sampling strategy, the influence of threshold variations, and the effect of adversarial suffix length are provided in Appendix 8.
Training dataset with prompts of similar semantics. We conduct a simple experiment by generating 50 normal prompts with almost the same semantics by GPT4 as the dataset $\mathcal{P}$. These prompts (listed in the Appendix 15) are generated from the prompt “Provide three pieces of advice for maintaining good health.”. For the target response, we use the “threatening” type. The suffix generated with our POUGH method only achieves 5% ASR on the test dataset, reflecting the importance of constructing a training dataset with high semantic diversity across prompts. Note that in this setting, we randomly select a prompt and calculate the semantic similarity between other prompts, achieving an average of 0.79. Also, for the diverse training set $\mathcal{P}$ sampled from $\mathcal{BP}$, randomly selecting a prompt and calculating the semantic similarity between other prompts, achieving an average of 0.31. The similarity observation reflects that the semantics extraction method (last hidden state of LLM) and similarity metric (cosine) used by us are effective.
Figure 2: Ablation study of ranking strategy.
TABLE IV: Ablation on size of dataset $\mathcal{P}$.
Size of Dataset $\mathcal{P}$
10 20 30 40 50 60 70 80
ASR (%) $\uparrow$ 46.4 75.0 76.1 86.6 92.6 85.7 90.7 88.2
Time (#NC) $\downarrow$ 1429 3067 2717 1548 2092 4875 7914 6064
6 Conclusion
We proposed POUGH, an efficient method for universal goal hijacking that combines an optimization algorithm with semantics-guided prompt organization. Our approach achieves high attack success rates while greatly reducing computational overhead. Unlike prior work focused solely on optimization, we emphasize the critical role of prompt diversity and ordering in improving universality and efficiency. In future work, we plan to explore more refined semantic similarity metrics.
7 Limitation
As an early work, we acknowledge that the proposed prompt organization strategies still have room for improvement. For example, maybe there exists more sophisticated semantic similarity metrics or improved methods for prompt sampling and ranking. However, although the prompts selected and sequence ranked by our strategies may not be the best choice, we firmly believe that our exploration is essential to emphasize the importance of prompt organization and serves as a valuable starting point for prompt-related research in the universal goal hijacking task.
Furthermore, given the limited research on cross-prompt universality, our study specifically focuses on this issue, without addressing cross-model universality. It is important to clarify that the primary objective of our work is to conduct an in-depth investigation into this particular problem and propose effective solutions, rather than attempting to encompass all possible aspects within a single study.